# The efficiency of black hole formation via collisions in stellar systems:

### v1

@@ Abstract
Context. This paper explores the theoretical relation between star clusters and black holes within, focusing on the potential role of Nuclear Star Clusters (NSCs), Globular Clusters (GCs), and Ultra Compact Dwarf Galaxies (UCDs) as environments that lead to black hole formation through stellar collisions.
Aims. The study aims to identify optimal conditions for stellar collisions in different stellar systems leading to the formation of very massive stars that subsequently collapse into black holes. Data from numerical simulations and observations of diverse stellar systems are analyzed, encompassing various initial conditions, initial mass functions, and stellar evolution scenarios.
Methods. We compute a critical mass, determined by the interplay of collision time, system age, and initial properties of the star cluster. The efficiency of black hole formation ($\epsilon_{\mathrm{BH}}$) is defined as the ratio of initial stellar mass divided by critical mass.
Results. The study finds out that stellar systems with a ratio of initial stellar mass over critical mass above 1 exhibit high efficiencies of black hole formation, ranging from $30-100\%$. While there is some scatter, potentially attributed to complex system histories and the presence of gas, the results highlight the potential for achieving high efficiencies through a purely collisional channel in black hole formation.
Conclusions. In conclusion, this theoretical exploration elucidates the connection between star clusters and black hole formation. The study underscores the significance of UCDs, GCs, and NSCs as environments conducive to stellar collisions leading to black hole formation. The defined black hole formation efficiency ($\epsilon_{\mathrm{BH}}$) is shown to be influenced by the ratio of initial stellar mass to critical mass.
@@ 1 Introduction
The origin of one of the most compact massive objects in the Universe, the supermassive black holes (SMBHs) has not been determined yet. These objects have masses ranging between $10^{6}-10^{10}\leavevmode\nobreak\ \mathrm{M_{\odot}}$ (Natarajan & Treister, 2009; Volonteri, 2010; Kormendy et al., 2010; Kormendy & Ho, 2013; McConnell et al., 2011; Zuo et al., 2015; King, 2016; Pacucci et al., 2017; Mehrgan et al., 2019; Onken et al., 2020). Some remarkable findings about SMBHs in the distant universe have been made by different research teams. The first $3$ quasars observed at redshift ($z$) larger than 6 were reported by Fan et al. (2003). Afterward Mortlock et al. (2011) observed a quasar at $z=7.085$ with a mass of $2\times 10^{9}\leavevmode\nobreak\ \mathrm{M}_{\odot}$. Soon after, Wu et al. (2015) discovered one of the most massive supermassive black holes known to date, weighing in at $1.2\times 10^{10}\leavevmode\nobreak\ \mathrm{M_{\odot}}$ at $z=6.3$. A few years later, Jeram et al. (2020) found a quasar at $z=2.89$ with a mass of $2.7\times 10^{10}\leavevmode\nobreak\ \mathrm{M_{\odot}}$, it is believed to be the most massive SMBH. Additionally, Wang et al. (2021) contributed to the exploration by spotting a distant Active Galactic Nucleus (AGN) at $z=7.642$, with a mass of $(1.6\pm 0.4)\times 10^{9}\leavevmode\nobreak\ \mathrm{M_{\odot}}$. Bañados et al. (2016) provided the observation of more than $100$ quasars with redshifts between $5.6\lesssim z\lesssim 6.7$. Lately, Fan et al. (2023) presented a review of high redshift quasars that include more than $300$ quasars at $z>6$.
@@
Some scenarios depict that SMBH formation is related to the remnants of Population III stars (Machida & Doi, 2013; Latif et al., 2014; Riaz et al., 2022). These stars usually form in metal-free clouds and accumulate mass onto their cores (Omukai & Nishi, 1998; Volonteri et al., 2003; Tan & McKee, 2004; Ricarte & Natarajan, 2018). Another possible route is the Direct Collapse model, where massive gas clouds collapse within atomic cooling halos at high redshift (Bromm & Loeb, 2003; Volonteri et al., 2008; Latif et al., 2013; Latif & Schleicher, 2015; Zwick et al., 2023) and runaway collisions and mergers of stars within star clusters (Rees, 1984; Portegies Zwart & McMillan, 2002; Fujii & Portegies Zwart, 2013; Katz et al., 2015; Sakurai et al., 2017; Reinoso et al., 2018, 2020; Sakurai et al., 2019; Vergara et al., 2021, 2023). Besides some authors proposed runaway collisions and mergers, with gas accretion (Boekholt et al., 2018; Regan et al., 2020; Chon & Omukai, 2020; Das et al., 2021; Schleicher et al., 2022, 2023; Reinoso et al., 2023). Runaway stellar collisions can lead to the formation of a Very Massive Star (VMS), which eventually results in a supernova and/or forms a stellar mass black hole (Heger & Woosley, 2002; Whalen et al., 2013; Chen et al., 2014). The interaction of the black hole with stars or other stellar mass black holes, besides their capacity to accrete surrounding material, can lead to the formation of an intermediate-mass black hole (IMBH) (Arca Sedda et al., 2023a, b, c). Stellar systems where this process may potentially occur include NSCs (Georgiev et al., 2016; Neumayer et al., 2020), GCs (Lützgendorf et al., 2011, 2012, 2013b), and UCDs (Seth et al., 2014; Ahn et al., 2017, 2018; Voggel et al., 2018). Regardless of their different ages, sizes, and masses (i.e. density), all of them offer suitable environments where stellar dynamics can unfold.
@@
NSCs are one of the densest stellar configurations known in the Universe, which usually fall in a mass range of $10^{5}-10^{7}\leavevmode\nobreak\ \mathrm{M_{\odot}}$ (Böker et al., 2002; Walcher et al., 2005; Côté et al., 2006; Böker, 2010; Georgiev et al., 2009, 2016; Fritz et al., 2016; Ordenes-Briceño et al., 2018; Sánchez-Janssen et al., 2019; Neumayer et al., 2020), with a half-mass radius of a few parsecs to a few tens of parsecs (Georgiev et al., 2016). Typically NSCs contain multiple stellar populations since they can host old stars but also present star formation (Walcher et al., 2005; Rossa et al., 2006; Kacharov et al., 2018). They can coexist with SMBHs (Neumayer & Walcher, 2012; Georgiev et al., 2016; Neumayer et al., 2020; Escala, 2021; Vergara et al., 2023). There are correlations of the SMBH mass with the mass of the galactic bulge (Magorrian et al., 1998; Marconi & Hunt, 2003; Häring & Rix, 2004), the SMBH mass and the velocity dispersion of the surrounding stars (Ferrarese & Merritt, 2000; Tremaine et al., 2002; Gültekin et al., 2009), as well as between the luminosity of the galaxy bulge and the NSC mass (Ferrarese et al., 2006; Rossa et al., 2006; Wehner & Harris, 2006; Böker, 2008; Savorgnan et al., 2016). Investigations of stellar orbits surrounding the SMBH at the center of the Milky Way (Ghez et al., 2008; Genzel et al., 2010; Gillessen et al., 2017), as well as the observation of the shadow of Sagittarius A* (Event Horizon Telescope Collaboration et al., 2022) provide evidence that there is a central massive object (Genzel et al., 2010; Schödel et al., 2014). The simultaneous presence of SMBHs and NSCs occur in galaxies with masses around $\sim 10^{10}\leavevmode\nobreak\ \mathrm{\mathrm{M_{\odot}}}$ (Seth et al., 2008). In less massive galaxies, that have less massive SMBHs, direct observations of their dynamical imprint on the NSCs are a challenging task that requires higher spatial resolution than for more massive galaxies at the same distance. Some studies propose the absence of a black hole or set an upper limit on its mass with $M_{\rm BH}<1500-3000\leavevmode\nobreak\ \mathrm{\mathrm{M_{\odot}}}$ (Gebhardt et al., 2001; Merritt et al., 2001). Other examples of NSCs with upper limits for black holes are NGC 300 and NGC 428 (Neumayer & Walcher, 2012). The gravitational influence of the Galactic center has a strong effect on the surrounding gas and stars, causing them to fall to the Galactic center by gradually losing their angular momentum (Shlosman et al., 1990; Escala, 2006). At high redshifts, the absence of SMBHs at Galactic centers and the abundance of gas lead to even stronger inflows, resulting in the formation of highly dense regions composed of stars and gas (Prieto & Escala, 2016). Other proposed accretion mechanisms include the clump migration through dynamical friction (Tremaine et al., 1975; Escala, 2007; Elmegreen et al., 2008; Arca-Sedda & Capuzzo-Dolcetta, 2014; Aharon & Perets, 2015) and gravitational torques during galaxy mergers (Barnes & Hernquist, 1991; Barnes, 2002; Mayer et al., 2010; Newton & Kay, 2013; Wurster & Thacker, 2013; Blumenthal & Barnes, 2018; Prieto et al., 2021).
@@
UCDs have a half-mass radius of approximately $\sim 10-50\leavevmode\nobreak\ \mathrm{pc}$ (Seth et al., 2014; Ahn et al., 2018; Faifer et al., 2017) but with masses of $10^{6}-10^{8}\leavevmode\nobreak\ \mathrm{M_{\odot}}$ (Dabringhausen et al., 2009, 2010), making them the densest stellar systems of the Universe. UCDs are old systems, but determining their age and metallicity can be challenging (Chilingarian & Mamon, 2008; Francis et al., 2012). Their origins are still unclear, with some theories proposing them to be massive star clusters formed through the merger of smaller GCs (Kroupa, 1998; Fellhauer & Kroupa, 2002; Fellhauer et al., 2002; Urrutia Zapata et al., 2019) or remnants of tidally stripped dwarf galaxies (Bassino et al., 1994; Bekki & Freeman, 2003; Pfeffer & Baumgardt, 2013) Some examples of these objects are UCD 330, UCD 320 (Voggel et al., 2018), M60-UCD1 (Seth et al., 2014), M59c0 (Ahn et al., 2017), M59-UCD-3 (Ahn et al., 2018), and FUCD3 (Afanasiev et al., 2018).
@@
GCs are stellar systems that have lower central density compared to NSCs with half-mass radii of a few to several dozen parsecs and masses of $10^{4}-10^{6}\leavevmode\nobreak\ \mathrm{M_{\odot}}$. These clusters are predominantly populated by primordial stars, signifying ages over 10 billion years (Kravtsov & Gnedin, 2005). They are found in the halos of galaxies, typically at distances of tens of thousands of light-years from the Galactic center. Most of the GCs in the Milky Way originate from a dual process, star formation during the early stages of cosmic evolution and the gradual disruption of dwarf galaxies (White & Rees, 1978; Kruijssen et al., 2019). Some examples of GCs are $\mathrm{\omega}$ Cen, 47Tuc, M15 or M80 (Harris, 1996). These stellar systems (NSCs, UCDs and GCs) provide an ideal environment for stellar collisions to occur.
@@
In general clusters with total mass below $10^{5}\leavevmode\nobreak\ \mathrm{M_{\odot}}$, which are not dense enough can form black holes with relatively low mass, typically amounting to a few percent of the initial mass (Portegies Zwart et al., 1999; Fujii & Portegies Zwart, 2013; Koliopanos, 2017; Arca Sedda et al., 2023a, b, c). Simulations of more massive clusters ($<10^{7}\leavevmode\nobreak\ \mathrm{M_{\odot}}$) are computationally expensive. However, adopting approximations, they can be explored via Fokker-Planck models at higher cluster masses (Lee, 1987; Quinlan & Shapiro, 1990; Stone et al., 2017). Escala (2021) proposed that the formation of massive objects in dense stellar systems is regulated by the relation between the collision time in a given system with the age and the relaxation time of that system, as it determines both whether collisions may efficiently occur and whether the system as a whole will be able to readjust to possible changes as a result of collisions. Vergara et al. (2023) performed N-body simulations to probe this scenario further. From the comparison of the collision time with the age of the system, they derived a critical mass $M_{\mathrm{crit}}$, this is the mass at which the collision timescale becomes equal to the age of the system. This implies that the large majority of stars in the system could then be expected to go through a collision, implying a drastic growth of the most massive objects. Indeed, Vergara et al. (2023) found that the mass of the most massive object systematically increases with the ratio of the initial mass divided by the critical mass ($M_{\mathrm{i}}/M_{\mathrm{crit}}$). Besides a very similar relation has been found in a preliminary comparison with observed NSCs.
@@
In this paper, we aim for a more systematic investigation of the same relation presented in Vergara et al. (2023), presenting and comparing simulation results from different investigations in terms of the ratio $M_{\mathrm{i}}/M_{\mathrm{crit}}$ and the growth of the most massive object. Particularly, the set of simulations analyzed here includes different radial profiles, initial mass functions as well as models for stellar evolution, e.g. employing simulations such as dragon II (Arca Sedda et al., 2023a, b, c), which are realistic simulations of GCs.
We further also refine the comparison with observational data, including additional NSCs, improved age estimates as well as additional stellar systems such as GCs and UCDs.
@@
The paper is organised as follows. Section 2 summarizes the methodology and data employed here. Section 3 shows the results of our analysis. Section 4 provides a summary of our conclusions drawn from this investigation.

### v2

# Efficiency of black hole formation via collisions in stellar systems

@@ Abstract
Context. This paper explores the theoretical relation between star clusters and black holes within them, focusing on the potential role of nuclear star clusters (NSCs), globular clusters (GCs), and ultra-compact dwarf galaxies (UCDs) as environments that allow for black hole formation via stellar collisions.
Aims. This study aims to identify the optimal conditions for stellar collisions across a range of stellar systems, leading to the formation of very massive stars that subsequently collapse into black holes. We analyze data from numerical simulations and observations of diverse stellar systems, encompassing various initial conditions, initial mass functions, and evolution scenarios.
Methods. We computed a critical mass, determined by the interplay of the collision time, system age, and initial properties of the star cluster. The efficiency of black hole formation ($\epsilon_{\mathrm{BH}}$) is defined as the ratio of initial stellar mass divided by the critical mass.
Results. We find that stellar systems with a ratio of initial stellar mass over critical mass above 1 exhibit a high efficiency in terms of of black hole formation, ranging from $30-100\%$. While there is some scatter, potentially attributed to complex system histories and the presence of gas, the results highlight the potential for achieving high efficiencies via a purely collisional channel in black hole formation.
Conclusions. In conclusion, this theoretical exploration elucidates the connection between star clusters and black hole formation. The study underscores the significance of UCDs, GCs, and NSCs as environments conducive to the black hole formation scenario via stellar collisions. The defined black hole formation efficiency ($\epsilon_{\mathrm{BH}}$) is shown to be influenced by the ratio of the initial stellar mass to the critical mass.
@@ 1 Introduction
The origin of one of the most compact massive objects in the Universe, namely, supermassive black holes (SMBHs) has not been determined thus far. These objects exhibit masses ranging between $10^{6}-10^{10}\leavevmode\nobreak\ \mathrm{M_{\odot}}$ (Natarajan & Treister, 2009; Volonteri, 2010; Kormendy et al., 2010; Kormendy & Ho, 2013; McConnell et al., 2011; Zuo et al., 2015; King, 2016; Pacucci et al., 2017; Mehrgan et al., 2019; Onken et al., 2020). Remarkable findings on SMBHs in the distant universe have been published by different research teams, starting from the first three quasars observed at redshift ($z$) higher than 6 reported by Fan et al. (2003). Afterward, Mortlock et al. (2011) observed a quasar at $z=7.085$ with a mass of $2\times 10^{9}\leavevmode\nobreak\ \mathrm{M}_{\odot}$. Soon after, Wu et al. (2015) discovered one of the most massive SMBHs known to date, weighing in at $1.2\times 10^{10}\leavevmode\nobreak\ \mathrm{M_{\odot}}$ at $z=6.3$. A few years later, Jeram et al. (2020) found a quasar at $z=2.89$ with a mass of $2.7\times 10^{10}\leavevmode\nobreak\ \mathrm{M_{\odot}}$, which is believed to be the most massive SMBH known thus far. Additionally, Wang et al. (2021) contributed to the exploration by spotting a distant active galactic nucleus (AGN) at $z=7.642$, with a mass of $(1.6\pm 0.4)\times 10^{9}\leavevmode\nobreak\ \mathrm{M_{\odot}}$. Bañados et al. (2016) provided the observation of more than $100$ quasars with redshifts between $5.6\lesssim z\lesssim 6.7$. Most recently, Fan et al. (2023) presented a review of high redshift quasars that includes more than $300$ quasars at $z>6$.
@@
Some scenarios depict SMBH formation as being related to the remnants of population III stars (Machida & Doi, 2013; Latif et al., 2014; Riaz et al., 2022). These stars usually form in metal-free clouds and accumulate mass onto their cores (Omukai & Nishi, 1998; Volonteri et al., 2003; Tan & McKee, 2004; Ricarte & Natarajan, 2018). Another possible route is the direct collapse model, where massive gas clouds collapse within atomic cooling halos at high redshift (Bromm & Loeb, 2003; Volonteri et al., 2008; Latif et al., 2013; Latif & Schleicher, 2015; Zwick et al., 2023) and runaway collisions and mergers of stars take place within star clusters (Rees, 1984; Portegies Zwart & McMillan, 2002; Fujii & Portegies Zwart, 2013; Katz et al., 2015; Sakurai et al., 2017; Reinoso et al., 2018, 2020; Sakurai et al., 2019; Vergara et al., 2021, 2023). In addition, some authors have proposed runaway collisions and mergers with gas accretion to explain their origins (Boekholt et al., 2018; Regan et al., 2020; Chon & Omukai, 2020; Das et al., 2021; Schleicher et al., 2022, 2023; Reinoso et al., 2023). Runaway stellar collisions can lead to the formation of a very massive star (VMS), which eventually results in a supernova and/or forms a stellar mass black hole (Heger & Woosley, 2002; Whalen et al., 2013; Chen et al., 2014). The interaction of the black hole with stars or other stellar mass black holes, besides their capacity to accrete surrounding material, can lead to the formation of an intermediate-mass black hole (IMBH) (Arca Sedda et al., 2023a, b, c). Stellar systems where this process may potentially occur include NSCs (Georgiev et al., 2016; Neumayer et al., 2020), GCs (Lützgendorf et al., 2011, 2012, 2013b), and UCDs (Seth et al., 2014; Ahn et al., 2017, 2018; Voggel et al., 2018). Regardless of their different ages, sizes, and masses (i.e., density), all of them offer suitable environments for stellar dynamics to unfold.
@@
As one of the densest stellar configurations known in the Universe, NSCs usually fall in the mass range of $10^{5}-10^{7}\leavevmode\nobreak\ \mathrm{M_{\odot}}$ (Böker et al., 2002; Walcher et al., 2005; Côté et al., 2006; Böker, 2010; Georgiev et al., 2009, 2016; Fritz et al., 2016; Ordenes-Briceño et al., 2018; Sánchez-Janssen et al., 2019; Neumayer et al., 2020), with a half-mass radius of a few parsecs to a few tens of parsecs (Georgiev et al., 2016). Typically, NSCs contain multiple stellar populations, since they can host old stars, however, they also present star formation (Walcher et al., 2005; Rossa et al., 2006; Kacharov et al., 2018). Also, they can coexist with SMBHs (Neumayer & Walcher, 2012; Georgiev et al., 2016; Neumayer et al., 2020; Escala, 2021; Vergara et al., 2023). There are correlations of the SMBH mass with the mass of the galactic bulge (Magorrian et al., 1998; Marconi & Hunt, 2003; Häring & Rix, 2004), the SMBH mass and the velocity dispersion of the surrounding stars (Ferrarese & Merritt, 2000; Tremaine et al., 2002; Gültekin et al., 2009), as well as between the luminosity of the galaxy bulge and the NSC mass (Ferrarese et al., 2006; Rossa et al., 2006; Wehner & Harris, 2006; Böker, 2008; Savorgnan et al., 2016). Investigations of stellar orbits surrounding the SMBH at the center of the Milky Way (Ghez et al., 2008; Genzel et al., 2010; Gillessen et al., 2017), as well as the observation of the shadow of Sagittarius A* (Event Horizon Telescope Collaboration et al., 2022) provide evidence that there is a central massive object (Genzel et al., 2010; Schödel et al., 2014). The simultaneous presence of SMBHs and NSCs occur in galaxies with masses around $\sim 10^{10}\leavevmode\nobreak\ \mathrm{\mathrm{M_{\odot}}}$ (Seth et al., 2008). In less massive galaxies, which have less massive SMBHs, direct observations of their dynamical imprint on the NSCs are a challenging task that requires higher spatial resolution than for more massive galaxies at the same distance. Some studies have proposed the absence of a black hole or have set an upper limit on its mass with $M_{\rm BH}<1500-3000\leavevmode\nobreak\ \mathrm{\mathrm{M_{\odot}}}$ (Gebhardt et al., 2001; Merritt et al., 2001). Other examples of NSCs with upper limits for black holes are NGC 300 and NGC 428 (Neumayer & Walcher, 2012). The gravitational influence of the Galactic center has a strong effect on the surrounding gas and stars, causing them to fall to the Galactic center by gradually losing their angular momentum (Shlosman et al., 1990; Escala, 2006). At high redshifts, the absence of SMBHs at Galactic centers and the abundance of gas lead to even stronger inflows, resulting in the formation of highly dense regions composed of stars and gas (Prieto & Escala, 2016). Other proposed accretion mechanisms include clump migration through dynamical friction (Tremaine et al., 1975; Escala, 2007; Elmegreen et al., 2008; Arca-Sedda & Capuzzo-Dolcetta, 2014; Aharon & Perets, 2015) and gravitational torques during galaxy mergers (Barnes & Hernquist, 1991; Barnes, 2002; Mayer et al., 2010; Newton & Kay, 2013; Wurster & Thacker, 2013; Blumenthal & Barnes, 2018; Prieto et al., 2021).
@@
With a half-mass radius of approximately $\sim 10-50\leavevmode\nobreak\ \mathrm{pc}$ (Seth et al., 2014; Ahn et al., 2018; Faifer et al., 2017) and masses of $10^{6}-10^{8}\leavevmode\nobreak\ \mathrm{M_{\odot}}$ (Dabringhausen et al., 2009, 2010), UCDs stand as the densest stellar systems of the Universe. They are known to be old systems, but determining their age and metallicity can be challenging (Chilingarian & Mamon, 2008; Francis et al., 2012). Their origins are still unclear, with some theories proposing them to be massive star clusters formed through the merger of smaller GCs (Kroupa, 1998; Fellhauer & Kroupa, 2002; Fellhauer et al., 2002; Urrutia Zapata et al., 2019) or remnants of tidally stripped dwarf galaxies (Bassino et al., 1994; Bekki & Freeman, 2003; Pfeffer & Baumgardt, 2013) Examples of these objects include UCD 330, UCD 320 (Voggel et al., 2018), M60-UCD1 (Seth et al., 2014), M59c0 (Ahn et al., 2017), M59-UCD-3 (Ahn et al., 2018), and FUCD3 (Afanasiev et al., 2018).
@@
Stellar systems with a lower central density compared to NSCs are known as GCs, with half-mass radii of a few to several dozen parsecs and masses of $10^{4}-10^{6}\leavevmode\nobreak\ \mathrm{M_{\odot}}$. These clusters are predominantly populated by primordial stars, indicating ages of over 10 billion years (Kravtsov & Gnedin, 2005). They are found in the halos of galaxies, typically at distances of tens of thousands of light years from the Galactic center. Most of the GCs in the Milky Way originate from a dual process: star formation during the early stages of cosmic evolution and the gradual disruption of dwarf galaxies (White & Rees, 1978; Kruijssen et al., 2019). Some examples of GCs are $\mathrm{\omega}$ Cen, 47Tuc, M15, or M80 (Harris, 1996). These stellar systems (NSCs, UCDs, and GCs) provide an ideal environment for stellar collisions to occur.
@@
In general, clusters with a total mass below $10^{5}\leavevmode\nobreak\ \mathrm{M_{\odot}}$ are not very dense, but can form black holes of a relatively low mass, typically amounting to a few percent of the initial mass (Portegies Zwart et al., 1999; Fujii & Portegies Zwart, 2013; Koliopanos, 2017; Arca Sedda et al., 2023a, b, c). Simulations of more massive clusters ($<10^{7}\leavevmode\nobreak\ \mathrm{M_{\odot}}$) are computationally expensive. However, by adopting approximations, they can be explored via Fokker-Planck models at higher cluster masses (Lee, 1987; Quinlan & Shapiro, 1990; Stone et al., 2017). Escala (2021) proposed that the formation of massive objects in dense stellar systems is regulated by the relation between the collision time in a given system with the age and the relaxation time of that system, as it determines both whether collisions may efficiently occur and whether the system as a whole will be able to readjust to possible changes as a result of collisions. Vergara et al. (2023) performed N-body simulations to probe this scenario further. From the comparison of the collision time with the age of the system, they derived a critical mass, $M_{\mathrm{crit}}$, which is the mass at which the collision timescale becomes equal to the age of the system. This implies that the large majority of stars in the system could then be expected to go through a collision, implying a drastic growth of the most massive objects. Indeed, Vergara et al. (2023) found that the mass of the most massive object systematically increases with the ratio of the initial mass divided by the critical mass ($M_{\mathrm{i}}/M_{\mathrm{crit}}$). In addition, a very similar relation has been found in a preliminary comparison with observed NSCs (Escala, 2021).
@@
In this paper, we aim for a more systematic investigation of the same relation reported in Vergara et al. (2023). We present and compare a range of simulation results from different investigations in terms of $M_{\mathrm{i}}/M_{\mathrm{crit}}$ values and the growth of the most massive object. In particular, the set of simulations analyzed here includes different radial profiles and initial mass functions, as well as models for stellar evolution; for instance, employing realistic simulations of GCs, such as dragon II (Arca Sedda et al., 2023a, b, c). We also aim to further refine the comparison with observational data, including additional NSCs, improved age estimates, and additional stellar systems such as GCs and UCDs.
@@
The paper is organised as follows. Section 2 summarizes the methodology and data employed. Section 3 shows the results of our analysis. Section 4 provides a summary of conclusions drawn from our investigation.